OpenStudy (anonymous):

2 x - 6 y = 5 x + y = 2 Solve the system of equations.

OpenStudy (anonymous):

substitution would be the easiest way to solve this... x + y = 2 x = -y + 2 now sub -y + 2 in for x in the other equation 2x - 6y = 5 2(-y + 2) - 6y = 5 can you finish ? solve for y, and once y is found, sub that value back into either equation and solve for x. Give it a try.

OpenStudy (anonymous):

yeah i can this now thx

OpenStudy (anonymous):

i cant find out plz tell me the answer